OpenStudy (australopithecus):

Two 150 g blocks, traveling on a frictionless surface, approach the opposite ends of a spring of force constant 1500 N/m. The blocks, traveling in opposite directions along the same straight line, have speeds of magnitude 1.2 m/s relative to the surface. What is the maximum compression of the spring in m when the blocks strike the spring?

OpenStudy (australopithecus):

I set up the equation W = Delta K -(1/2)(1500)x^2 = 0 - 2*(1/2)(0.15)(1.2)^2

OpenStudy (australopithecus):

which give me 0.01697 which I don't think is correct as these are the possible answers and I would think they would round up if I do not multiply the kinetic energy by 2 I get 0.012 but my reasoning to do so is because it is being compressed on both ends by the same kinetic energy a. 0.012 b. 0.013 c. 0.014 d. 0.015 e. 0.016

OpenStudy (australopithecus):

I wish these practice sheets had solutions on them how are we suppose to know we are solving the problem correctly if they dont even supply the answer ugh

OpenStudy (anonymous):

I am not seeing any problems in your work. If you made a mistake, I did also.

OpenStudy (australopithecus):

damn

OpenStudy (ujjwal):

I guess it should be :\[\frac{1}{2}K(2x)^2\]on LHS